User Tools

Site Tools


Tid 70m observed at 1726-1800 UT.

DAS Profile:

Note added on DOY 298: It appears the frequency set up for this experiment might have been wrong, recording 1613 MHz +/- 16MHz instead of 1668 MHz +/- 16 MHz. We will send the data anyway just in case.

lbaops/lbaoct2016/rk16bltilog.txt · Last modified: 2016/10/25 14:38 by shoriuchi